1. Find V. 10 kΩ 2VW- 50 kΩ ΕΛΛΑ 20 ΚΩ 3 V-ΛΑ- 50 ΚΩ -1V.-- 2. Find V. 10 Ω 5Ω + Λ 1V + + 12 Ω 3Ω
Answer #1

① Find out lok 50% mo w 20K vo 3 V ye Sok. -v. Apply KCL. 2-V2, = 7 3-12 20k +::1 - V2 į VX-VO 50 Jok 50k Yx=0 - 2 3 gok •102 Vou + 14 vo Valam 122 32 111 Va-1 V 1st opam Apply the Ker Vy-o 32 Væ-701 12-2 Voi 12 - - ادر) 12 VO! 4-1 12 12 Nor=-3Y A-Vo = Voix2 Vo = - Vox Z Vo =-(-3)/2 Vo=68